OpenStudy (anonymous):

what is the simplified form of x^4-81/x+3

OpenStudy (anonymous):

@Michele_Laino do u think u can help me by any chance?

OpenStudy (welshfella):

x^4 - 81 can be regarded as the difference of 2 squares (x^2)^2 - 9^2 - can you factor that? - remember that a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

how do i factor that? @welshfella

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

relace a by x^2 and b by 9 and see what you get

OpenStudy (anonymous):

can u put that in a formula for me then i will solve it best i can

OpenStudy (welshfella):

a^2 - b^2 = ( a + b)( a - b) just replace a by x^2 and b by 9

OpenStudy (anonymous):

so (x^2)^2-81?

OpenStudy (welshfella):

(x^2)^2 - 9^2 = ?

OpenStudy (anonymous):

(x^2+9)(x^2-9) ?? @welshfella

OpenStudy (welshfella):

right now do you think any os those 2 factors will factor any further?

OpenStudy (welshfella):

one of them will

OpenStudy (welshfella):

we have another 'difference of 2 squares'

OpenStudy (anonymous):

uhhhh idrk could u explain that more?

OpenStudy (welshfella):

x^2 - 9 = ?

OpenStudy (welshfella):

take the 2 square roots, giving:- x^2 - 9 = ( x - 3)(x + 3)

OpenStudy (welshfella):

right? - similar thing to the first step

OpenStudy (welshfella):

have a good look at it I think you'll see it i gotta go right now

OpenStudy (anonymous):

x^3-3x^2+9x-27 ?? @welshfella

OpenStudy (welshfella):

sorry must go

OpenStudy (anonymous):

:( was i right tho? @welshfella

OpenStudy (welshfella):

the x +3 cancesl out x^4 - 81 = (x^2 + 9)(x + 3)(x - 3) -------- -------------------- = (x^2 + 9(x - 3) x + 3 (x + 3)

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

so i was right? @welshfella

OpenStudy (welshfella):

yup

OpenStudy (anonymous):

thanks!

OpenStudy (welshfella):

yw